Find five consecutive numbers that the sum of the frist and the last is 84. Find the numbers.
Answer by stanbon(75887)

Find five consecutive numbers that the sum of the frist and the last is 84. Find the numbers.
-----
1st: x-2
2nd: x-1
3rd x
4th: x+1
5th: x+2
--------
Equation:
sum = 84
2x = 84
x = 42
-----
x-2 = 40
x-1 = 41
x = 42
x+1 = 43
x+2 = 44
-------------
Cheers,
Stan H.
=============